OpenStudy (anonymous):

rationalize the denominator and simplify squareroot of 10/ squareroot of 7

OpenStudy (anonymous):

answer is root70/7

OpenStudy (anonymous):

sqrt 10/sqrt[7] Multiply top and bottom by sqrt 7

OpenStudy (anonymous):

oops it was \[\frac{\sqrt{10}}{\sqrt{7}}\] my mistake

OpenStudy (anonymous):

correct ans is \[\sqrt{70} / 7\]